RC meeting and Run Plan, Saturday November 19th

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

No RC meeting on Saturday and Sunday

RC Updates

Summary of swing 11/18 - day 11/19

  • Continued production run on polarized He3 target @45uA beam current with prescale: ps1=9, ps2=15, ps3=0, ps5=3 and all other -1
  • DAQ crashed few times during swing and owl shift, restarted the platform (or xterms) to get it fixed
  • Bill worked around the oven temperature to make it more uniform https://logbooks.jlab.org/entry/4087154 , but it seems higher oscillation rate with lower temp variation
  • Recent NMR sweep target polarization 45.4% (at 11:50 am) and 45.8% (at 2:40pm) today

Run plan for until tomorrow

  • Continue production data taking on polarized He3
  1. 45uA of beam current | Use ps1=9, ps2=15, ps3=0, ps5=3, rest =-1 | One hour long runs
  2. Move the target to the pickup coil position, and do NMR sweep around every 3 hours
  3. Return to step 1

Issues

  • DAQ rate dropping during even if beam is on => Expert need to figure out issues and permanent solution
  • Dump right Ion chamber cabling needs to be changed => Jerry Kowal will try to fix for short term by changing cable on next longer downtime